# Characterizations of Pseudolinear and Semistrictly Quasilinear Functions

### v1

@@ Abstract
In this paper we obtain several new complete characterizations of pseudolinear functions. Two of the results are of first-order and one is derivative free. All results are derived in terms of the Clarke-Rockafellar subdifferential. Additionally, we prove a characterization of the semistrictly quasilinear functions. It is similar to the derivative free characterization of the pseudolinear functions. We also find the conditions such that a semistrictly quasilinear function become pseudolinear.
Keywords: Pseudolinear functions, semistrictly quasiliear functions, nonsmooth analysis, conditions for pseudoliearity and semistrictly quasiliearity
@@ 1 Introduction
The concepts of pseudolinearity and semistrict quasilinearity provide in a very natural way generalizations of the linearity. These classes of functions include many useful functions. Among them are, for instance, the linear fractional functions. Several interesting properties of these functions appeared in [7, 3, 8, 6].
In the paper by Ivanov [1] were obtained some new characterizations of pseudoconvex functions and semistrictly ones. Here we apply the obtained conditions for pseudoconvexity to derive three complete characterizations of pseudolinear functions in terms of the Clarke’s generalized directional derivative (Theorems 2, 3, 5). Theorem 2 is a generalization to non-differentiable functions of a theorem due to Chew and Choo [3]. In Theorem 4, we obtain necessary conditions for pseudolinearity, but they are not sufficient for pseudolinearity. It is interesting which is the largest class of functions such that these conditions hold. We prove that they become necessary and sufficient when the function is semistrictly quasilinear, which implies that the mentioned class is the exact one. Every pseudolinear function, which is subdifferentiable with respect to the Clarke-Rockafellar subdifferential, is semistrictly quasilinear. We find the conditions that a semistrictly quasilinear function have to satisfy to become pseudolinear.
2 Preliminaries
In this section, we present the necessary and sufficient conditions for a given function to be pseudoconvex, which were derived in [1]. In the next sections, we apply these results.
In the sequel, we suppose that $\mathbb{E}$ is a Banach space. We denote by $\mathbb{E}^{*}$ its dual and the duality pairing between the vectors $a\in\mathbb{E}^{*}$ and $b\in\mathbb{E}$ by $\langle a,b\rangle$, by $\mathbb{R}$ the set of reals.
@@ 2 Preliminaries
Let $f$ be a locally Lipschitz function, defined on some open set $X$ in a Banach space $\mathbb{E}$. The Clarke’s generalized derivative at the point $x\in X$ in direction $v\in\mathbb{E}$ is defined by
@@
The Clarke’s subdifferential (or Clarke’s generalized gradient) of $f$ at $x$ is defined as follows:
@@
Recall that a real function is said to be quasiconvex iff,
@@
The following result is due to Daniilidis, Hadjisavvas [5, Proposition 2.2].
@@
The next theorem were derived by Ivanov [1]:
Theorem 1.
@@ 3 Characterizations of pseudolinear functions
The following result is a particular case of Lemma 3 in the paper [1]:
Lemma 2.
Let $f$ be a locally Lipschitz pseudoconvex real function, defined on some open convex set in a Banach space $\mathbb{E}$, which contains the convex set $S$. Then the following implication holds
$x\in S,\;y\in S,\;f(y)\leq f(x)\quad\Rightarrow\quad\langle x^{*},y-x\rangle\leq 0,\quad\forall x^{*}\in\partial f(x).$
@@
Let $f$ be a locally Lipsctitz $f$ is pseudolinear on $S$ with respect to the Clarke’s directional derivative if and only if there exists a positive function $p:S\times S\times\mathbb{E}^{*}\to(0,+\infty)$ with
@@
Let $x\in S$, $y\in S$, $x^{*}\in\partial f(x)$ and equation (3) is satisfied. Obviously (4) implies that $f$ is pseudolinear. ∎
Theorem 2 is a generalization to non-differentiable functions of a known claim due to Chew and Choo [3, Proposition 2.1].
@@
The converse claim follows from Theorem 1. ∎
@@
$f[x+\lambda(y-x)]\leq\max\{f(x),f(y)\}=f(x)\quad{\rm forall)}\;\lambda\in[0,1],$
$f[x+\lambda(y-x)]\geq\min\{f(x),f(y)\}=f(x)\quad{\rm forall)}\;\lambda\in[0,1].$
@@
(b) there is a function $b:S\times S\times[0,1]\to(0,+\infty)$ such that for all $x\in S$, $y\in S$ there exists the limit $\lim_{\lambda\downarrow 0}b(x,y,\lambda), q(x,y)q(x,y)$ is strictly positive, and for each $\lambda\in[0,1]$ equation (8) and inequality (9) are satisfied.
@@
we consider several cases:
First, $f(y)<f(x)$. Then $f(y)<f(z(\lambda))$ for all sufficiently small $\lambda>0$. It follows from $q=1/p$ and (4) that
@@
Second, $f(y)>f(x)$. We have $f(y)>f(z(\lambda))$ for all sufficiently small $\lambda>0$. Thanks to (4) and (15), we obtain that (16) is satisfied again, and (14) holds.
Third, $f(y)=f(x)$. We have that $\nabla f(x)(y-x)=0$. Therefore $q(x,y)=1$. It this case $f[x+\lambda(y-x)]=f(x)$ for all $\lambda\in[0,1]$, and $b(x,y,\lambda)=1$ for every $\lambda\in[0,1]$. Therefore, the required equality (14) is satisfied. It is seen from (10) that for all $x$ and $y$ such that $f(y)\neq f(x)$ we have $q(x,y)=\langle\nabla f(x),y-x\rangle/[f(y)-f(x)]>0$.
@@
$f(x+\lambda(y-x))-f(x)]/\lambda=(x_{1}y_{2}-x_{2}y_{1})/[x_{1}(x_{1}+\lambda(y_{1}-x_{1})]$
@@
$b(x,y,\lambda)=(f(x+\lambda(y-x)))/(\lambda(f(y)-f(x)))=y_{1}/(x_{1}+\lambda(y_{1}-x_{1})).$
@@ 4 On semistrictly quasilinear and pseudolinear functions
The claim follows directly from the definition of semistrict quasiconvexity and Theorem 4. ∎
@@
The case $f(y)>f(x)$ is similar. It follows from semistrict quasiconvexity that $f[x+\lambda(y-x)]<f(y)$. Therefore $f[x+\lambda(y-x)]-f(x)<[f(y)-f(x)]$, which implies that (17) is also satisfied. Therefore, Condition (8) holds and $0<\lambda b<1$.
Let $x\in S$, $y\in S$, $f(x)=f(y)$. By Proposition 1 ff is both quasiconvex and quasiconcave. Therefore,
@@
Let $S$ be an convex set in a Banach space $E$, $f$ be a Fréchet differentiable semistrictly quasiconvex and semistrictly quasiconcave function, defined on some open set $\Gamma$, containing $S$. Then $f$ is pseudolinear on $S$ if and only if the following implication holds:
@@
Let $S$ be a convex subset of an open set $\Gamma$ in a Banach space $\mathbb{E}$. Suppose that $f:\Gamma\to\mathbb{R}$ is a continuous function, which is both semistrictly quasiconvex and semistrictly quasiconcave on $S$ and $\partial^{\uparrow}f(x)\neq 0$, $\partial^{\uparrow}(-f)(x)\neq 0$ for all $x\in S$. Then $f$ is pseudolinear with respect to the Clarke-Rockafellar subdifferential if and only if the both implications hold:
@@
Let $f$ be pseudolinear. We prove implication (19). Take arbitrary $x\in S$, $y\in S$, $\xi\in\partial f(x)$ such that $\langle\xi,y-x\rangle=0$. If $f(y)<f(x)$, then by pseudoconvexity we have $\langle\xi,y-x\rangle<0$, which is a contradiction. The proof of implication (20) is similar.

### v2

@@ Abstract
In this paper, we obtain several new complete characterizations of pseudolinear functions. Two of the results are of first-order and one is derivative free. The results are derived in terms of the Clarke-Rockafellar subdifferential. Additionally, we prove a characterization of the semistrictly quasilinear functions. It is similar to the derivative free characterization of the pseudolinear functions. We also find the conditions such that a semistrictly quasilinear function become pseudolinear.
Keywords: Pseudolinear functions, semistrictly quasiliear functions, nonsmooth analysis, conditions for pseudoliearity and semistrict quasiliearity
@@ 1 Introduction
The concepts of pseudolinearity and semistrict quasilinearity provide in a very natural way generalizations of the linearity. These classes include many useful functions. Among them are, for instance, the linear fractional functions. Several interesting properties of these of pseudolinear functions appeared in [6, 14, 15, 21]. More generalized convex functions have been studied for example in the books [5, 9, 18, 19, 20]. Quasilinear functions have been studied in [19] under the name quasimonotone functions.
In this paper, we obtain more characterizations of pseudolinear functions. In Theorem 3, we obtain necessary conditions for pseudolinearity, but they are not sufficient. The respective necessary and sufficient conditions are obtained in Theorem 4. Theorem 1 is well known and we prove it for reader’s convenience. Really it is a generalization to non-differentiable functions of a theorem due to Chew and Choo [6]. It is interesting which is the largest class of functions such that the conditions of Theorem 3 hold. We prove that they become necessary and sufficient when the function is semistrictly quasilinear, which implies that the mentioned class is the exact one. Every pseudolinear function, which is subdifferentiable with respect to the Clarke-Rockafellar subdifferential, is semistrictly quasilinear. We also find the conditions such that a semistrictly quasilinear function have to satisfy to become pseudolinear.
In the sequel, we suppose that $\mathbb{E}$ is a Banach space. We denote by $\mathbb{E}^{*}$ its dual and the duality pairing between the vectors $a\in\mathbb{E}^{*}$ and $b\in\mathbb{E}$ by $\langle a,b\rangle$, by $\mathbb{R}$ the set of reals, by $\overline{\mathbb{R}}$ the union $\mathbb{R}\cup\{+\infty\}$, by $B(x,r)$ the closed ball of a center $x$ with a radius $r$. Let $f:X\to\mathbb{R}$ be a proper extended real-valued function, whose domain is the set
${\rm dom}(f):=\{x\in\mathbb{E}\mid f(x)<+\infty\}.$
@@
Let $f:X\to\mathbb{R}$ be a proper extended real-valued function and $x\in{\rm dom}(f)$. The Clarke-Rockafellar generalized derivative of $f$ at $x$ in direction $v$ is defined by
$f^{\uparrow}(x,v)=\sup_{\varepsilon>0}\,\limsup_{(y,\alpha)\downarrow{}_{f}x;t\downarrow 0}\,\inf_{u\in B(v,\varepsilon)}\,[f(y+tu)-\alpha]/t,$
where $(y,\alpha)\downarrow{}_{f}x$ means that $y\to x$, $\alpha\to f(x)$, $\alpha\geq f(y)$ (see [23]), and $y\to x$ implies that the norm $\Arrowvert y-x\Arrowvert$ approaches 0. If $f$ happens to be lower semicontinuous at $x$ the definition can be expressed in the slightly simpler form
$f^{\uparrow}(x,v)=\sup_{\varepsilon>0}\,\limsup_{y\downarrow{}_{f}x;t\downarrow 0}\,\inf_{u\in B(v,\varepsilon)}\,[f(y+tu)-f(y)]/t,$
where $y\downarrow{}_{f}x$ means that $y\to x$, $f(y)\to f(x)$. When $f$ is locally Lipschitz, this derivative coincides with the Clarke generalized derivative [7], which is defined by
@@
The Clarke-Rockafellar subdifferential of $f$ at $x$ is defined as follows:
$\partial^{\uparrow}f(x)=\{x^{*}\in\mathbb{E}^{*}\mid\langle x^{*},v\rangle\leq f^{\uparrow}(x,v),\quad\forall v\in\mathbb{E}\}$
with the convention that $\partial^{\uparrow}f(x)=\emptyset$ if $x\notin{\rm dom}(f)$. The Clarke’s subdifferential (or Clarke’s generalized gradient) of $f$ at $x$ is defined as follows:
@@
Recall that a real function is said to be quasiconvex on a convex set $X$ iff,
@@
The following result is due to Daniilidis, Hadjisavvas [8, Proposition 2.2].
@@
The following result is a particular case of Lemma 3 in the paper by Ivanov [10]:
Lemma 2.
Let $f$ be a lower semicontinuous proper extended pseudoconvex real function, defined on some open convex set in a Banach space $\mathbb{E}$, which contains the convex set $S$. Then the following implication holds
$x\in S,\;y\in S,\;f(y)\leq f(x)\quad\Rightarrow\quad\langle x^{*},y-x\rangle\leq 0,\quad\forall x^{*}\in\partial^{\uparrow}f(x).$
The next result were also derived by Ivanov [10]:
Lemma 3.
@@ 2 Characterizations of pseudolinear functions
The next theorem is well known. Its proof appears in terms of various derivatives and subdifferential in [1, 6, 17, 22]. It was proved initially in [6]. Our proof is more general, but similar. We prove it, because the construction in it is used in the proof of another result.
@@
Let $f$ be a lower semicontinuous proper extended function with a convex domain $S$. Then $f$ is pseudolinear on $S$ with respect to the Clarke-Rockafellar directional derivative if and only if there exists a positive function $p:S\times S\times\mathbb{E}^{*}\to(0,+\infty)$ with
@@
Let $x\in S$, $y\in S$, $x^{*}\in\partial^{\uparrow}f(x)$ and equation (3) is satisfied. Obviously (3) implies that $f$ is pseudolinear. ∎
@@
The converse claim follows from Lemma 3. ∎
Equation (5) is a type of monotonicity of the subdifferential of a pseudolinear function. More properties of such maps are studied in [4]. More types monotone maps have been defined in [12, 13].
@@
$f[x+\lambda(y-x)]\leq\max\{f(x),f(y)\}=f(x)\quad{\rm forall}\;\lambda\in[0,1],$
$f[x+\lambda(y-x)]\geq\min\{f(x),f(y)\}=f(x)\quad{\rm forall}\;\lambda\in[0,1].$
@@
(b) there is a function $b:S\times S\times[0,1]\to(0,+\infty)$ such that for all $x\in S$, $y\in S$ there exists the limit
$q(x,y)=\lim_{\lambda\downarrow 0}b(x,y,\lambda),$ (13)
$q(x,y)$ is strictly positive, and for each $\lambda\in[0,1]$ equation (8) and inequality (9) are satisfied.
@@
we consider two cases:
First, $f(y)\neq f(x)$. Then $f(y)\neq f(z(\lambda))$ for all sufficiently small $\lambda>0$. It follows from $q=1/p$ and (4) that
@@
Second, $f(y)=f(x)$. We have that $\nabla f(x)(y-x)=0$. Therefore $q(x,y)=1$. It this case $f[x+\lambda(y-x)]=f(x)$ for all $\lambda\in[0,1]$, and $b(x,y,\lambda)=1$ for every $\lambda\in[0,1]$. Therefore, the required equality (15) is satisfied. It is seen from (4) that for all $x$ and $y$ such that $f(y)\neq f(x)$ we have $q(x,y)=\langle\nabla f(x),y-x\rangle/[f(y)-f(x)]>0$.
@@
$(f(x+\lambda(y-x))-f(x))/\lambda=(x_{1}y_{2}-x_{2}y_{1})/(x_{1}(x_{1}+\lambda(y_{1}-x_{1})))$
@@
$b(x,y,\lambda)=(f(x+\lambda(y-x))-f(x))/(\lambda(f(y)-f(x)))=y_{1}/(x_{1}+\lambda(y_{1}-x_{1})).$
@@ 3 On semistrictly quasilinear and pseudolinear functions
The claim follows directly from the known theorem that every pseudoconvex function is semistrictly quasiconvex. ∎
@@
The case $f(y)>f(x)$ is similar. It follows from semistrict quasiconvexity that $f[x+\lambda(y-x)]<f(y)$. Therefore $f[x+\lambda(y-x)]-f(x)<f(y)-f(x)$, which implies that (18) is also satisfied. Therefore, Condition (8) holds and $0<\lambda b<1$.
Let $x\in S$, $y\in S$, $f(x)=f(y)$. By Proposition 1 $f$ is both quasiconvex and quasiconcave. Therefore,
@@
The following claim is well known [15].
Proposition 3.
Let $S$ be an open convex set in a finite-dimensional space $E$, $f$ be a Fréchet differentiable function, defined on $S$. Then $f$ is pseudolinear on $S$ if and only if the following sets are equal for all $x\in S$: $\{y\in S:\;\nabla f(x)(y-x)=0\}$ and $\{y\in S:\;f(y)=f(x)\}$.
The next theorem is similar, but different from Proposition 3.
@@
Let $S$ be an convex set in a Banach space $E$, $f$ be a Fréchet differentiable semistrictly quasilinear function, defined on some open set $\Gamma$, containing $S$. Then $f$ is pseudolinear on $S$ if and only if the following implication holds:
@@
Let $S$ be a convex subset of an open set $\Gamma$ in a Banach space $\mathbb{E}$. Suppose that $f:\Gamma\to\mathbb{R}$ is a continuous function, which is both semistrictly quasiconvex and semistrictly quasiconcave on $S$ and $\partial^{\uparrow}f(x)\neq 0$, $\partial^{\uparrow}(-f)(x)\neq 0$ for all $x\in S$. Then $f$ is pseudolinear with respect to the Clarke-Rockafellar subdifferential if and only if both implications hold:
@@
Let $f$ be pseudolinear. We prove implication (20). Take arbitrary $x\in S$, $y\in S$, $\xi\in\partial f(x)$ such that $\langle\xi,y-x\rangle=0$. If $f(y)<f(x)$, then by pseudoconvexity we have $\langle\xi,y-x\rangle<0$, which is a contradiction. we have $\langle\xi,y-x\rangle>0$, which is also a contradiction. Therefore $f(y)=f(x)$. Thus, implication (20) is proved. The proof of implication (21) is similar.